### v2

# On non-isomorphic universal sofic groups

Abstract.
We show that there are $2^{\aleph_{0}}$ non-isomorphic universal sofic groups. This proves a conjecture of Simon Thomas.
1. Introduction
Starting with the work of Gromov [MR1694588] and Weiss [MR1803462], there has been growing interest in the class of sofic groups over the last two decades. Elek and Szabó [MR2178069] were the first to characterize countable sofic groups by the property of embedability into a metric ultraproduct of symmetric groups with respect to the normalized Hamming metric; following their work these metric ultraproducts were termed universal sofic groups. See Pestov’s excellent survey for more background [MR2460675]. Metric ultraproducts of finite groups and related constructions have been a subject of intense study eversince, see [MR2298607, MR2607888, MR3710760, MR3846318, MR3210125, MR3162821, MR4292938, MR4232187, MR4141378, MR3966829, MR3749196, MR3299505]. It is elementary to see that the embedability of countable sofic groups does not depend on the particular choice of metric ultraproduct, i.e., on the choice of the ultrafilter. Thus, universal sofic groups cannot be distinguished by their countable subgroups. However, Thomas clarified in [MR2607888] that assuming the negation of the continuum hypothesis, there must be $2^{2^{\aleph_{0}}}$ non-isomorphic universal sofic groups and naturally raised the question, how the situation unfolds under the assumption of the continuum hypothesis.
Our main result, which proves a conjecture by Thomas [MR2607888, Conjecture 1.2], is the following theorem:
Theorem 1.1.
There are at least $2^{\aleph_{0}}$ pairwise non-isomorphic metric ultraproducts $\prod^{\rm met}_{\mathcal{U}}({\rm Sym}(n),d_{n})$.
The result is independent of the question whether we want to assume the continuum hypothesis or not. However, if we assume the continuum hypothesis, it follows from classical arguments that the bound $2^{\aleph_{0}}$ is tight. Phrasing our result in slightly different language, we prove that the FO-theory in the sense of model theory for metric structures (see [benyaacov]) of the symmetric group $({\rm Sym}(n),d_{n})$ endowed with the normalized Hamming metric
$d_{n}(\sigma,\tau):=\frac{|\{1\leq i\leq n|\sigma(i)\neq\tau(i)\}|}{n}$
does not converge as $n$ tends to infinity. This answers [MR2607888, Question 1.3] and raises the natural question of the lowest complexity of the FO-sentence (in the sense of model theory for metric structures), that is able to distinguish two sequences of symmetric groups.
Our proof of Theorem 1.1 uses a blend of recent results revolving around groups with Kazhdan’s property (T), see [kazhdan-T], starting with recent breakthrough results of Caprace–Kassabov [caprace2023tame] and Bartholdi–Kassabov [bartholdi2023property] in providing Kazhdan groups that surject onto a sufficiently large set of alternating groups. We use Kun’s analysis of sofic approximation of Kazhdan groups, see [MR4555893, kun2021expanders], and the application of these results to the study of almost centralizers of sofic approximations of Kazhdan groups obtained by Kun and the second author in [kun2019inapproximability]. Finally, we rely on recent work of Becker and Chapman [MR4634678], who answered a question from [kun2019inapproximability] and proved stability of uniform almost homomorphisms from finite groups to permutation groups. Finally, we use more classical work of Felgner [MR1107758] and Wilson [MR1477188] in giving characterizations of finite non-abelian simple groups in terms of first order sentences in the language of group theory.
Let’s outline the argument: The basic mechanism of the proof is to observe that the metric groups $({\rm Sym}(n!\!),d_{n!})$ behave in different ways in the sense of model theory for metric structures for different $n\in\mathbb{N}.$ This is due to the existence of the left-right action of ${\rm Sym}(n)\times{\rm Sym}(n)$ on the set ${\rm Sym}(n)$, where the factors are centralizers of each other and each of them acts freely. In the presence of sufficient expansion, this situation is very rigid and can be uniquely recovered in a suitable metric ultraproduct $\prod^{\rm met}_{\mathcal{U}}({\rm Sym}(n!\!),d_{n!})$. Thus, the algebraic ultraproduct $\prod^{\rm alg}_{\mathcal{U}}{\rm Sym}(n)$ can be defined inside the metric ultraproduct $\prod^{\rm met}_{\mathcal{U}}({\rm Sym}(n!\!),d_{n!})$ as the centralizer of a particular finite set of elements. In order to be able to make uniform statements, we need the construction of Bartholdi–Kassabov mentioned above. Now, it is folklore how to distingish algebraic ultraproducts of symmetric groups using first order sentences in the language of group theory and this finishes the proof. The actual work and novelty of the arguments lies in the effort to prove the required rigidity of centralizers and double centralizers. As explained above, this makes crucial use of particular expansion generators of ${\rm Sym}(n)$, rigidity results of Kun and the second author, and stability of uniform almost homomorphisms.
2. Almost homomorphisms
2.1. Local almost homomorphisms and sofic approximations
We will be very brief on the definition of a sofic group and refer to [MR2460675] for details, examples, and references.
Definition 2.1.
Let $\Gamma$ be a group. Let $F\subset\Gamma$ be finite and $\delta>0$. We say that a map $\sigma\colon\Gamma\to{\rm Sym}(n)$ is a $(F,\delta)$-homomorphism if $d_{n}(\sigma(gh),\sigma(g)\sigma(h))\leq\delta$ for all $g,h\in F$. We say that $\sigma$ is $(F,\delta)$-injective if $d_{n}(\sigma(g),\sigma(h))\geq 1-\delta$ for all $g,h\in F$ with $g\neq h.$
Definition 2.2.
A group $\Gamma$ is sofic if and only if for any finite set $F\subset\Gamma$ and $\delta>0$, $\Gamma$ admits an $(F,\delta)$-injective $(F,\delta)$-homomorphism to ${\rm Sym}(n)$ for some $n$.
In case of countable groups, which is the most interesting case, it is easy to see that we can characterize soficity by requiring the existence of a sequence of maps $(\sigma_{n})_{n}$ that are $(F,\delta)$-injective $(F,\delta)$-homomorphisms for larger and larger $F$ and smaller and smaller $\delta>0.$
Let $\mathcal{U}$ be a non-principal ultrafilter on $\mathbb{N}$. We consider the algebraic ultraproduct
$\prod{\vphantom{\prod}}^{\mathrm{alg}}_{\mathcal{U}}{\rm Sym}(n)\coloneqq\left(\prod_{n}{\rm Sym}(n)\right)/\left\{(\sigma_{n})_{n}\in\prod_{n}{\rm Sym}(n)\,\Big{|}\,\sigma_{n}=1_{n}\text{ along }\mathcal{U}\right\}$
and define the universal sofic group with respect to the ultrafilter $\mathcal{U}$ to be
$\operatorname{Sym}(\mathcal{U})\coloneqq\prod\vphantom{\prod}^{\mathrm{met}}_{\mathcal{U}}({\rm Sym}(n),d_{n})=\left(\prod_{n}{\rm Sym}(n)\right)/{\rm N}(\mathcal{U}),$
where
${\rm N}(\mathcal{U})=\left\{(\sigma_{n})_{n}\in\prod_{n}{\rm Sym}(n)\,\Big{|}\,\lim_{n\to\mathcal{U}}d_{n}(1_{n},\sigma_{n})=0\right\}.$
Note that ${\rm Sym}(\mathcal{U})$ carries a natural bi-invariant metric $d_{\mathcal{U}}$ arising as the ultralimit of the metrics on ${\rm Sym}(n).$ This metric takes values in the interval $[0,1].$
Definition 2.3.
A subgroup $\Gamma\leq{\rm Sym}(\mathcal{U})$ is said to be $1$-discrete if $d_{\mathcal{U}}(g,h)=1$ for all $g,h\in\Gamma$ with $g\neq h.$
The following result goes back to [MR2178069], where universal sofic groups were considered for the first time and many basic results about them were proved.
Theorem 2.4 (Elek–Szabó).
Let $\Gamma$ be a countable group. The following conditions are equivalent.
  1. $(i)$
    The group $\Gamma$ is sofic.
  2. $(ii)$
    The group $\Gamma$ is isomorphic to a $1$-discrete subgroup of ${\rm Sym}(\mathcal{U})$ for all non-principal ultrafilters $\mathcal{U}.$
  3. $(iii)$
    The group $\Gamma$ is isomorphic to a subgroup of ${\rm Sym}(\mathcal{U})$ for some non-principal ultrafilter $\mathcal{U}.$
The basic observation that leads to the previous theorem is that a sequence of maps $\sigma_{n}\colon\Gamma\to{\rm Sym}(m_{n})$, where $\sigma_{n}$ is a $(F_{n},\delta_{n})$-injective $(F_{n},\delta_{n})$-homomorphism, with $\lim_{n\to\infty}\delta_{n}=0$ and $(F_{n})_{n}$ a sequence of finite subsets that increases to $\Gamma$, leads naturally to an injective homomorphism
$\sigma=[\sigma_{n}]_{n}\colon\Gamma\to{\rm Sym}(\mathcal{U})$
with $\mathcal{U}$ an ultrafilter supported on the set $\{m_{n}\mid n\in\mathbb{N}\}$. Moreover, the image is clearly $1$-discrete. Conversely, any such $1$-discrete embedding arises from a sequence as above. We say that two sequences $(\sigma_{n})_{n}$ and $(\sigma^{\prime}_{n})_{n}$ are essentially the same (along equivalent ultrafilters $\mathcal{U},\mathcal{V}$, see 5.3 for details) if they lead to the same embedding, i.e., $\sigma=\sigma^{\prime}.$
As a particular consequence of Theorem 2.4, we note again that existence of countable subgroups cannot distinguish the groups ${\rm Sym}(\mathcal{U})$ for different ultrafilters up to isomorphism.
If $\Gamma$ is finitely generated, there is a slightly different picture of sofic approximations as follows: Let $S\subset\Gamma$ be a finite generating set. Any $(F,\delta)$-injective $(F,\delta)$-homomorphism $\sigma\colon\Gamma\to{\rm Sym}(n)$ can be restricted to $S$ and leads to a Schreier graph ${\rm Sch}(n,\sigma(S))$ of the free group on $S$. It is sometimes more convenient to work with sequences of Schreier graphs that look locally more and more like the Cayley graph ${\rm Cay}(\Gamma,S)$ instead of families of maps as in 2.2. We will frequently switch between the different pictures.
2.2. Uniform almost homomorphisms
A different notion of almost homomorphism arises when we require the almost multiplicativity to occur in a uniform way on the entire domain of the map.
Definition 2.5.
Let $G$ be a group and $\delta>0$. We say that a map $\sigma\colon G\to{\rm Sym}(n)$ is a uniform $\delta$-homomorphism if
$d_{n}(\sigma(gh),\sigma(h)\sigma(h))\leq\delta,\quad\forall g,h\in G.$
Following standard terminology, we say that two maps $\sigma,\sigma^{\prime}\colon G\to{\rm Sym}(n)$ are uniformly $\varepsilon$-close, if $d_{n}(\sigma(g),\sigma^{\prime}(g))\leq\varepsilon$ for all $g\in G.$ Uniform perturbations of homomorphisms are the trivial examples of uniform almost homomorphisms. Uniform almost homomorphisms were first studied by Kazhdan [kazhdan] in a more functional analytic setting. The theme was further developed in [burgerozawathom, hatamigowers, dechiffreozawathom]. The setting of uniform almost homomorphisms with permutation group targets comes up naturally in the study of almost centralizers of sofic approximations of Kazhdan groups, see [kun2019inapproximability]. The most striking rigidity/stability result for uniform almost homomorphisms in the permutation group setting is the following theorem, which answered a question from [kun2019inapproximability]:
Theorem 2.6 (Becker–Chapman, [MR4634678]*Theorem 1.2).
For every $\varepsilon>0$, there exists $\delta>0$, such that for every $n\in\mathbb{N}$, every finite group $G$ and every uniform $\delta$-homomorphism $\sigma\colon G\to{\rm Sym}(n)$, there exists a homomorphism $\pi\colon G\to{\rm Sym}(m)$ with $m\in[n,(1+\varepsilon)n]$ which is uniformly $\varepsilon$-close to $\sigma\oplus 1_{m-n}.$
In fact the proof shows that $\delta=\varepsilon/2039$ is enough.
3. Applications of Kazhdan’s property (T)
3.1. Kazhdan groups and their finite quotients
Already in the work of Thomas [MR2607888], the existence of expander generators of symmetric groups played a crucial role. There, the original work of Kassabov [kassabov2007] was used in order to obtain a certain rigidity of the centralizer of a certain finite set of elements in a universal sofic group. In our approach, we need to use more refined result in the same direction. Indeed, in the meantime, a question of Lubotzky was answered positively, and the existence of a Kazhdan group with infinitely many alternating quotients was proven. We will need the following consequence of [bartholdi2023property]*Corollary 14.
Theorem 3.1 (Bartholdi–Kassabov).
There exists a Kazhdan group $\Gamma$, which surjects onto ${\rm Alt}(p^{4}-1)$ for all primes $p\geq 13.$
It would be more convenient for the proof of our main theorem to have a Kazhdan group $\Gamma$, which surjects onto all alternating groups in a way such that any non-trivial $g\in\Gamma$ becomes trivialized only in finitely many of these quotients. However, it seems to be out of reach to construct such a group.
3.2. Sofic approximations of Kazhdan groups
The work of Gábor Kun [kun2021expanders], anwering a question of Bowen, was the first to prove a positive structure result for sofic approximation beyond the amenable case.
Theorem 3.2 (Kun).
Every sofic approximation of a Kazhdan group is essentially a disjoint union of expanders.
In subsequent work of Kun and the second author [kun2019inapproximability] it was realized that sofic approximations of Kazhdan groups $\Gamma$ admit a special rigidity when it comes to the study of almost automorphisms of such. Let $S$ be a finite generating set for $\Gamma$. For a given $(F,\delta)$-injective $(F,\delta)$-homomorphism $\sigma\colon\Gamma\to{\rm Sym}(n)$ (where without loss of generality $F\supseteq S$) it makes sense to study the set of $\varepsilon$-automorphisms. These are bijections $\rho\in{\rm Sym}(n)$ of the vertex set of the associated Schreier graph ${\rm Sch}(n,\sigma(S))$, such that
|$|\{(\rho(i),s,\rho(j))\in E\mid(i,s,j)\in E\}|\geq(1-\varepsilon)|E|,$
i.e., the map $\rho$ is almost a graph automorphism. If ${\rm Sch}(n,\sigma(S))$ is a $c$-expander, then it follows that the set of $\varepsilon$-automorphisms of ${\rm Sch}(n,\sigma(S))$ consists of clusters, i.e., any two $\varepsilon$-automorphisms are either $O_{c}(1/n)$-close or $(1-O_{c}(1/n))$-apart (see [kun2019inapproximability] for the details). It thus makes sense to talk about the set of clusters. Naturally, one would expect that the set of clusters forms a group. However, the evident problem is that the product of two $\varepsilon$-homomorphisms is a priori only a $2\varepsilon$-homomorphism, so that we run out of our set of clusters unless some form of self-improvement mechanism can be implemented. One of the main technical results in [kun2019inapproximability] is that this is indeed possible in the situation just discussed.
Theorem 3.3 ([kun2019inapproximability, see Theorem 3.4 and Lemma 4.2]).
Let $\Gamma$ be a Kazhdan group with a finite generating set $S\subset\Gamma$, $c>0$, and $\varepsilon>0$ be sufficiently small. There exists a finite set $F\subset\Gamma$ and $\delta>0$, such that for any $(F,\delta)$-injective $(F,\delta)$-homomorphism $\sigma\colon\Gamma\to{\rm Sym}(n)$ such that: if ${\rm Sch}(n,\sigma(S))$ is a $c$-expander, then the set of clusters $G$ of $\varepsilon$-automorphisms of ${\rm Sch}(n,\sigma(S))$ is a group.
We call the group $G$ obtained from the previous theorem the cluster group of $\varepsilon$-automorhisms or the $\varepsilon$-centralizer of $\sigma\colon\Gamma\to{\rm Sym}(n)$. If $G$ is the cluster group of $\varepsilon$-automorphisms as above, we may choose a representative of each cluster and obtain a uniform $C\varepsilon$-homomorphism $\sigma\colon G\to{\rm Sym}(n)$ for some $C$ depending only on the generating set and Kazhdan constant of $\Gamma$. This observation will be used in combination with Theorem 2.6 in the proof of the main theorem.
We also obtain an immediate corollary.
Corollary 3.4.
The centralizer of a sofic approximation of a Kazhdan group by expanders is an algebraic ultraproduct of finite groups. More precisely, let $\Gamma$ be a Kazhdan group, $S\subset\Gamma$ be a finite generating set, and $\sigma=[\sigma_{n}]_{n}\colon\Gamma\to\prod_{\mathcal{V}}^{\rm met}({\rm Sym}(n),d_{n})$ is a sofic approximation by expanders, and $\varepsilon>0$ sufficiently small. Then the centralizer of $\sigma(\Gamma)$ is $1$-discrete and isomorphic to the algebraic ultraproduct $\prod_{\mathcal{V}}G_{n}$ of finite groups $G_{n}$, where $G_{n}$ is the $\varepsilon$-centralizer of $\sigma_{n}(S)$.
Note that the assumption that the sofic approximation is by expanders is crucial for the $1$-discreteness of the centralizer. As it turns out, there is a partial converse to this result: $1$-discreteness of the centralizer and the double centralizer guarantees that the approximation is by expanders.
Lemma 3.5.
Let $\sigma\colon\Gamma\to\prod_{\mathcal{V}}^{\rm met}({\rm Sym}(n),d_{n})$ be a sofic approximation of an infinite Kazhdan group. If the centralizer $C(\sigma(\Gamma))$ and the double centralizer $C(C(\sigma(\Gamma)))$ are $1$-discrete, then the sofic approximation is essentially by expanders, i.e., the Schreier graphs obtained from the sofic approximation are asymptotically equivalent to a $c$-expander sequence, where $c>0$ only depends on the Kazhdan constant of $\Gamma$.
Proof.
By Theorem 3.2, we can assume that the graphs $X_{n}$ induced by our sofic approximation of $\Gamma$ are disjoint unions of expanders; let $Y_{n}$ be a component of maximal size in $X_{n}$.
We claim that $\lim\limits_{n\to\mathcal{U}}|Y_{n}|/|X_{n}|=1$. Indeed, suppose that $\lim\limits_{n\to\mathcal{U}}|Y_{n}|/|X_{n}|=\sigma<1$. If $\sigma>0$, then the sofic approximation decomposes as the direct product $\sigma=\sigma_{1}\times\sigma_{1}^{c}$ of its restrictions to $Y_{n}$ and $Y_{n}^{c}$, and therefore the centralizer $C(\sigma(\Gamma))$ contains $C(\sigma_{1}(\Gamma))\times C(\sigma_{2}(\Gamma))$. If $C(\sigma_{1}(\Gamma))$ is non-trivial, the elements of the form $(c_{1},1)\in C(\sigma_{1}(\Gamma))\times C(\sigma_{2}(\Gamma))$ have support at most $\alpha<1$ which contradicts the assumption that $C(\sigma(\Gamma))$ is $1$-discrete. If $C(\sigma_{1}(\Gamma))$ is trivial, then $C(C(\sigma_{1}(\Gamma)))$ contains the ultraproduct of $\operatorname{Sym}(Y_{n})$ which contradicts the assumption that $C(C(\sigma(\Gamma)))$ is $1$-discrete. The case $\alpha=0$ is similar. Indeed, now all components are small and we can replace $Y_{n}$ by a disjoint union of components of size approximately $|X_{n}|/2$ and repeat the argument above. This finishes the proof. ∎
4. The first order theory of finite groups
We do not claim any novelty for the results in this section. However, we found it convenient to recall some results in detail and write our own proofs of some basic observations.
4.1. Characterizing alternating groups
The following result is due to Felgner [MR1107758], see the work of Wilson [MR1477188, Theorem 5.1] for a proof.
Theorem 4.1 (Felgner).
There exists a FO-sentence $\varphi$ in the language of groups which is satisfied by a finite group if and only if it is non-abelian simple.
For convenience, let us spell out a concrete FO-sentence that characterizes simplicity among non-abelian finite groups. Following Wilson’s account of Felgner’s result, we set:
$\varphi_{1}=\left(\forall g\ \forall h\ (g\neq 1\wedge C(g,h)\neq\{1\})\rightarrow\left(\cap_{k}(C(g,h)C(C(g,h)))^{k}=\{1\}\right)\right)$
$\varphi_{2}=(\forall g\ \exists h_{1},h_{2}\ (g=[h_{1},h_{2}])).$
Then, a finite group satisfies $\varphi\coloneqq\varphi_{1}\wedge\varphi_{2}$ if and only if it is non-abelian simple. Starting with this FO-sentence, it is possible to characterize more specific families of non-abelian finite simple groups:
Proposition 4.2.
There exists a FO-sentence in the language of groups that is satisfied for a simple group if and only if it is alternating.
Sketch of proof:.
By the classification of finite simple groups, we may ignore sporadic groups and focus on the classical families. We may also ignore abelian groups and groups of fixed bounded rank. It remains to distuingish ${\rm Alt}(n)$ among the other families of finite simple groups of Lie type of unbounded rank, untwisted or twisted. If such a finite simple group (of rank high enough) is defined over a field of characteristic $p$, then the behaviour of centralizers of elements of order $q$ prime to $p$ is different from the behaviour of the centralizer of a non-trivial unipotent element of order $p$. Indeed, the centralizer of an element of order $q$ is a central product of quasi-simple groups whereas the centralizer of a non-trivial unipotent element has a potentially large abelian normal subgroup. This allows to recognise the characteristic of the underlying field using the FO-properties of the non-abelian simple group. Moreover, it is easy to see that alternating groups show the exceptional behaviour for all primes. Whence, alternating groups are the only non-abelian simple groups among the remaining cases which show the exceptional behaviour for two different primes. Thus, we can characterize alternating groups among finite simple groups using a suitable constructed FO-sentence. This finishes the proof. ∎
Proposition 4.3.
Let $q\geq 3$ be a prime number and $l\in\mathbb{Z}$ with $10\leq l\leq q-$ 1 . There exists a FO-sentence $\varphi(l,q)$ in the language of groups such that a finite group $G$ satisfies $\varphi(l,q)$ if and only if $G$ is isomorphic to ${\rm Alt}(n)$ with $n\equiv l\mod q.$
Proof.
By 4.2, we may already assume that $G$ is an alternating group. After replacing $l$ by $q+l$ if necessary, we may assume that $l\geq 4$. We set
$\varphi(l,q)=(\exists g\ (g^{q}=1)\wedge[C(g):{\rm Alt}(l)\times C(g,{\rm Alt}(l))]\leq 2),$
i.e., the sentence says that there exists an element $g$ of order $q$ such that the centralizer of $g$ has a subgroup of index $2$, which is a product of groups, one factor being isomorphic to ${\rm Alt}(l)$ and the other factor being its centralizer in $C(g)$. It is elementary to check that this happens only when $g$ is the union of a number of $q$-cycles and exactly $l$ fixed points. Note that we can define the centralizer of $g$, require existence of a copy of ${\rm Alt}(l)$ for fixed $l$, define its centralizer and their product in FO-language. Also the statement that those groups are contained in each other with index $2$ is expressible in FO-language. This finishes the proof. ∎
Results like the previous proposition are well-known, see for example the proof of Theorem 3.1 in [thomasetal]. It is natural to wonder what kind of subsets of the natural numbers can be defined in this way. More precisely:
Question 4.4.
For which subsets $A\subset\mathbb{N}$ does there exist a FO-sentence $\varphi$ in the language of groups such that $n\in A$ if and only if ${\rm Alt}(n)$ satisfies $\varphi$?
Remark 4.5.
As an example, note that the sentence
$\exists g\ \forall h\ (gh=hg)\rightarrow(h=1\vee(\exists k\ hk=kg))$
characterizes ${\rm Sym}(n)$ with $n=p$ or $n=p+1$ and $p$ a prime within the set of all symmetric groups. Hence, more exotic sets than mere arithmetic progressions are definable using FO-sentences in the language of groups.
4.2. Non-isomorphic ultraproducts of alternating groups
The following lemma is an elementary consequence of Dirichlet’s theorem on primes in arithmetic progressions. We denote the set of prime numbers by $\mathbb{P}$.
Lemma 4.6.
For every prime $q\geq 7$, there exist residues $a_{q,0},a_{q,1}\in\{0,\dots,q-2\}$ with $a_{q,0}\neq a_{q,1}$ such that the following holds: For any finite set of prime numbers $Q\subset\mathbb{P}_{\geq 7}$ and any choice $\gamma\colon Q\to\{0,1\}$, there exists a prime $p\geq 13$, such that $p^{4}-1$ is congruent to $a_{q,\gamma(q)}$ modulo $q$ for all $q\in Q.$
Proof.
For every $q\geq 7$, the multiplicative group of residues $(\mathbb{Z}/q\mathbb{Z})^{\times}$ is not $4$-torsion. Hence, there is a non-zero fourth power different from the residue class of $1$. We set $a_{q,0}=0$ and $a_{q,1}=c-1$, where $c$ is a representative of a non-trivial fourth power and $d^{4}\equiv c\mod q$. We also set $b_{q,0}=1$ and $b_{q,1}:=d$. Now, let $Q\subset\mathbb{P}_{\geq 7}$ be finite and $\gamma$ be as above. By the Chinese Remainder Theorem, there exists $l\in\mathbb{N}$ which is congruent to $b_{q,\gamma(q)}$ modulo $q$ for all $q\in Q.$ Note that ${\rm gcd}(l,\prod_{q\in Q}q)=1$. Now, by Dirichlet’s theorem, there exists a prime $p\geq 13$, which is congruent to $l$ modulo $\prod_{q\in Q}q.$ Clearly, $p$ is congruent to $b_{q,\gamma(q)}$ modulo $q$ for all $q\in Q$. However, this implies that $p^{4}-1$ is congruent to $a_{q,\gamma(q)}$ for all $q\in Q$, as required. ∎
Theorem 4.7.
There are $2^{\aleph_{0}}$ ultrafilters on $\mathbb{N}$ supported on the set $\{p^{4}-1\mid p\geq 13,p\mbox{ prime}\}$ such that the ultraproducts $\prod^{\mathrm{alg}}_{\mathcal{U}}\operatorname{Alt}(m)$ are pairwise non-isomorphic.
Proof.
For any function $\gamma\colon\mathbb{P}_{\geq 7}\to\{0,1\}$, we can construct an ultrafilter $\mathcal{U}(\gamma)$ such that
(1) $\{p^{4}-1\in\mathbb{N}\mid p\in\mathbb{P}_{\geq 13},p^{4}-1\equiv_{q}a_{q,\gamma(q)}\}\in\mathcal{U}(\gamma),\quad\forall q\in\mathbb{P}_{\geq 7}.$
Indeed, for any finite part of $\gamma$, there exists a prime $p\geq 13$ solving the problem by the previous lemma. We define an ultrafilter $\mathcal{U}(\gamma)$ by choice of a limit point of the net of those primes as the finite part increases.
Now, since FO-sentences can distinguish congruence conditions by 4.3, we conclude that the resulting groups $\prod^{\rm alg}_{\mathcal{U}(\gamma)}\operatorname{Alt}(m)$ are pairwise non-elementarily equivalent. Indeed, by the observation above, the group $\prod^{\rm alg}_{\mathcal{U}(\gamma)}\operatorname{Alt}(m)$ satisfies the FO-sentences $\varphi(a_{q,\gamma(q)},q)$ and $\neg\varphi(a_{q,1-\gamma(q)},q)$ for all primes $q\geq 7.$
The ultrafilters $\mathcal{U}(\gamma)$ will be crucial in the proof of the main theorem. We will denote by
$\mathbb{U}\coloneqq\{\mathcal{U}(\gamma)\mid\gamma\colon\mathbb{P}_{\geq 7}\to\{0,1\}\}\subset\beta\mathbb{N}\setminus\mathbb{N}$
the set of ultrafilters constructed in the previous theorem.
5. The main theorem
5.1. Isomorphisms of metric ultraproducts
Recall that there is a natural metric on a metric ultraproduct of symmetric groups, arising as the ultralimit of the normalized Hamming metrics. The following proposition is folklore:
Proposition 5.1.
Every isomorphism of universal sofic groups is automatically isometric.
Proof.
We need to encode the metric using the group theoretic properties of ${\rm Sym}(\mathcal{U})$. For $g\in{\rm Sym}(\mathcal{U})$, we denote by ${\rm cl}(g)$ its conjugacy class. Consider
$\Sigma:=\{g\in{\rm Sym}(\mathcal{U})\mid\forall k\geq 1\ g\in{\rm cl}(g^{k})\}.$
One can show that the set $\Sigma$ consists of all those elements in the ultraproduct that can be represented by permutations with just one non-trivial cycle and a number of fixed points.
Let $r\in\mathbb{Q}$ and consider the set $B(r):=\{g\in{\rm Sym}(\mathcal{U})\mid d_{\mathcal{U}}(1,g)<r\},$ i.e., the open ball of radius $1/n$ around the identity element. We claim that
$\Sigma\cap B(1/n)=\{g\in\Sigma\mid({\rm cl}(g))^{n}\neq{\rm Sym}(\mathcal{U})\}$
and
$\Sigma\cap B(m/n)=(\Sigma\cap B(1/n))^{m}.$
This shows that any isomorphism must preserve the sets $\Sigma\cap B(r)$ for $r\in\mathbb{Q}.$
Now, the centralizer of each element $g\in{\rm Sym}(\mathcal{U})$ decomposes uniquely as a product $C(g)=S_{\infty}\times S_{1}\times\prod_{k\geq 2}(A_{k}\rtimes S_{k})$, where each $S_{k}$, for $k\geq 1$, is a (potentially trivial) metric ultraproduct of symmetric groups and $A_{k}$ is the group of $(\mathbb{Z}/k\mathbb{Z})$-valued functions (modulo measure zero) on the corresponding Loeb space. The factors $S_{k}$, for $k\geq 1$, in the decomposition corresponds to the product of all $k$-cycles in the decomposition of $g$, whereas $S_{\infty}$ corresponds to the part, where $g$ acts with larger and larger cycles. It is easy to see, that the size $\lambda_{k}\in[0,1]$ of the support of each of the subgroups $S_{k}$ is measured by the support of the largest element in $\Sigma$ that lies in the subgroup. In particular, by our reasoning above, the sizes are preserved by any isomorphism. Since
$d_{\mathcal{U}}(1,g)=\sum_{k\neq 1}\lambda_{k}=1-\lambda_{1},$
we conclude that the metric is preserved. ∎
One can compare the previous result with related results obtained by Păunescu, see [MR3299505].
Definition 5.2.
We say that two non-principal ultrafilters $\mathcal{U},\mathcal{V}$ on $\mathbb{N}$ are equivalent if there exists a monotone map $\kappa\colon\mathbb{N}\to\mathbb{N}$, such that $\kappa_{*}(\mathcal{U})=\mathcal{V}$ and $\lim_{n\to\mathcal{U}}\kappa(n)/n=1.$
It is elementary to check that equivalence of ultrafilters (in this sense) defines an equivalence relation on the set of ultrafilters. It is a well-known observation that universal sofic groups with respect to equivalent ultrafilters are isomorphic.
Proposition 5.3.
Let $\mathcal{U},\mathcal{V}$ be equivalent, non-principal ultrafilters on $\mathbb{N}$. Then, the two metric ultraproducts $\prod_{\mathcal{U}}^{\rm met}({\rm Sym}(n),d_{n})$ and $\prod_{\mathcal{V}}^{\rm met}({\rm Sym}(n),d_{n})$ are isomorphic.
Consider the map $f(n)=n!\!/2$. In the proof of the main theorem in the next section, ultrafilters of the form $f_{*}(\mathcal{U})$ will play an important role. We will rely on the following basic observation:
Lemma 5.4.
Let $\mathcal{U}\neq\mathcal{V}\in\beta\mathbb{N}$ be two ultrafilters. Then, $f_{*}(\mathcal{U})$ and $f_{*}(\mathcal{V})$ are not equivalent.
Proof.
Note that $f(n)/f(m)\not\in(1/2,2)$ if $n\neq m$. Suppose that $f_{*}(\mathcal{U})$ is equivalent to $f_{*}(\mathcal{V})$. Now, this implies that $\kappa(n)=n$ along $f_{*}(\mathcal{U})$. In particular, $f_{*}(\mathcal{U})=f_{*}(\mathcal{V})$ and hence $\mathcal{U}=\mathcal{V}$. ∎
5.2. Proof of the main theorem
Our aim is to show that the metric ultraproducts $\prod_{f_{*}(\mathcal{U})}^{\rm met}({\rm Sym}(n),d_{n})$ are pairwise non-isomorphic when $\mathcal{U}$ ranges among the ultrafilters constructed in Theorem 4.7. Then, Theorem 1.1 is an immediate consequence. In fact, we show a slightly stronger result:
Theorem 5.5.
Let $\mathcal{U}\in\mathbb{U}$ be an ultrafilter and $\mathcal{V}$ be any other ultrafilter. Assume that $\prod_{f_{*}(\mathcal{U})}^{\rm met}({\rm Sym}(n),d_{n})$ and $\prod_{\mathcal{V}}^{\rm met}({\rm Sym}(n),d_{n})$ are isomorphic. Then there exists a unique ultrafilter $\mathcal{U}^{\prime}\in\beta\mathbb{N}$ such that $\mathcal{V}$ is equivalent to $f_{*}(\mathcal{U}^{\prime}).$ Moreover, the ultraproducts $\prod^{\rm alg}_{\mathcal{U}}{\rm Alt}(n)$ and $\prod^{\rm alg}_{\mathcal{U}^{\prime}}{\rm Alt}(n)$ are isomorphic.
In particular, the metric ultraproducts $\prod_{f_{*}(\mathcal{U})}^{\rm met}({\rm Sym}(n),d_{n})$ with $\mathcal{U}\in\mathbb{U}$ are pairwise non-isomorphic.
The proof of Theorem 5.5 will be carried our in several steps combining the results prepared in the above sections.
Proof.
Using the result of Bartholdi–Kassabov (Theorem 3.1), we fix a Kazhdan group $\Gamma$ together with surjective homomorphisms $r_{p}\colon\Gamma\to\operatorname{Alt}(p^{4}-1)$ for every prime $p\geq 13$. We set $\Lambda\coloneqq(\mathbb{Z}/2\mathbb{Z})\ltimes(\Gamma\times\Gamma)$, where $\mathbb{Z}/2\mathbb{Z}$ acts by permuting the factors. We will call the corresponding subgroups the left resp. right copy of $\Gamma$ in $\Lambda$ and denote them by $\Gamma_{\ell}$ resp. $\Gamma_{r}$. Note that by combining the left-right action of $\Gamma$ on itself with the quotient map $r_{p}$ we get a sequence of homomorphisms $q_{p}\colon\Lambda\to\operatorname{Sym}((p^{4}-1)!\!/2)$, where $\mathbb{Z}/2\mathbb{Z}$ acts by inversion. For every $\mathcal{U}\in\mathbb{U}$, they combine to a homomorphism $\pi_{\mathcal{U}}\colon\Lambda\to\prod_{f_{*}(\mathcal{U})}^{\rm met}({\rm Sym}(n),d_{n}).$ We make the following claims:
  1. (1)
    The centralizer $C(\pi_{\mathcal{U}}(\Gamma_{l}))$ and the double centralizer $C(C(\pi_{\mathcal{U}}(\Gamma_{l})))=C(\pi_{\mathcal{U}}(\Gamma_{r}))$ are $1$-discrete, isomorphic to $\prod^{\rm alg}_{\mathcal{U}}{\rm Alt}(n),$ and interchanged by $\mathbb{Z}/2\mathbb{Z}$. This follows from 3.4 together with the fact, that any sequence of finite quotients of a property (T) group yields expander graphs, see [margulis-expanders].
  2. (2)
    The centralizer $C(\pi_{\mathcal{U}}(\Gamma_{\ell}))$ satisfies the FO-sentences described in Theorem 4.1 and 4.2.
Now, let’s assume that $\prod_{f_{*}(\mathcal{U})}^{\rm met}({\rm Sym}(n),d_{n})$ and $\prod_{\mathcal{V}}^{\rm met}({\rm Sym}(n),d_{n})$ are isomorphic. Therefore, we may consider $\pi_{\mathcal{U}}$ as a homomorphism from $\Lambda$ to the group $\prod_{\mathcal{V}}^{\rm met}({\rm Sym}(n),d_{n})$. Property (1) from above is transferred to $\prod_{\mathcal{V}}^{\rm met}({\rm Sym}(n),d_{n})$ by 5.1. Obviously, Property (2) is also transferred via the isomorphism. We note that the image $\Gamma_{\mathcal{U}}\coloneqq\pi_{\mathcal{U}}(\Gamma_{\ell})$ is an infinite Kazhdan group and $\pi_{\mathcal{U}}$ can naturally be considered to be a sofic approximation of this group. The aim of the proof is to recover the finitary situation completely along the ultrafilter $\mathcal{V}$.
First we observe that in view of 3.5 applied to $\Gamma_{\mathcal{U}}$, we can assume that our sofic approximation of $\Gamma_{\mathcal{U}}$ yields a $c$-expander sequence, where $c>0$ depends only on the Kazhdan constant for $\Gamma$. Therefore, taking a positive and small enough $\delta<c/10^{4}$, we can now apply 3.4 to $\Gamma_{\mathcal{U}}$. Since the finite groups $G_{n}$ satisfy the FO-sentences from (2), we deduce that $G_{n}\cong\operatorname{Alt}(m_{n})$ for some $m_{n}\in\mathbb{N}$. Similarly, the centralizer of $\pi_{\mathcal{U}}(\Gamma_{r})$ follows to be an algebraic ultraproduct of finite groups $G^{\prime}_{n}\cong\operatorname{Alt}(m^{\prime}_{n})$. Since the action of $\pi_{\mathcal{U}}(\mathbb{Z}/2\mathbb{Z})$ interchanges them, we see that $m_{n}=m^{\prime}_{n}$ along $\mathcal{U}$. Moreover, we can take the permutations $s_{n}\coloneqq\sigma_{n}(s)$, $s\in S$, as a subset of generators of $G_{n}$, and we fix lifts $t_{n}\in\operatorname{Sym}(n)$ of $\pi_{\mathcal{U}}(t)$ to $\operatorname{Sym}(n)$ of the generator $t\in\mathbb{Z}/2\mathbb{Z}$, so that conjugation by $t_{n}$ implements an isomorphism $G_{n}\cong G_{n}^{\prime}$.
We have now obtained a uniform $\delta$-homomorphism
$\sigma_{n}\colon(\mathbb{Z}/2\mathbb{Z})\ltimes(\operatorname{Alt}(m_{n})\times\operatorname{Alt}(m_{n}))\to\operatorname{Sym}(n).$
By Theorem 2.6, it is uniformly $2039\delta$-close to a homomorphism $\pi_{n}\colon(\mathbb{Z}/2\mathbb{Z})\ltimes(\operatorname{Alt}(m_{n})\times\operatorname{Alt}(m_{n}))\to\operatorname{Sym}(k)$ with $k\in[n,(1+2039\delta)n]$.
We claim that $\left|n-\frac{m_{n}!}{2}\right|\leq C\delta n$ for some universal $C>0$. Indeed, since $\pi_{n}(s_{n})$ are $2039\delta$-close to $\sigma_{n}(s)\oplus 1_{k-n}$ in the Hamming distance, the Schreier graphs of $\pi_{n}(G_{n})$ have a connected component of size $(1-2039\delta/c)n$. Therefore the action of $\pi_{n}(G_{n})$ is transitive on a subset $Y_{n}\subset\{1,\dots,k\}$ of size at least $(1-2039\delta/c)n$. Since the action of $\pi_{n}(G_{n}^{\prime})$ can only permute the connected components, it has to preserve this connected component. Altogether we get a transitive action $G_{n}\curvearrowright Y_{n}$ whose centralizer is isomorphic to $G_{n}$. This implies that the action $(G_{n}\curvearrowright Y_{n})$ is isomorphic to the action of $G_{n}$ on itself. Indeed, if $Y_{n}\cong G_{n}/H_{n}$ and $z\in\operatorname{Sym}(G_{n}/H_{n})$ commutes with the action of $G_{n}$, then $z$ is uniquely determined by $z(H_{n})\in G_{n}/H_{n}$, so that the size of the centralizer cannot exceed $|G_{n}/H_{n}|$. In particular, we get $|Y_{n}|=m_{n}!\!/2$, and the claim follows.
Now, we see that $\mathcal{V}$ is equivalent to $f_{*}(\mathcal{U})$ for an ultrafilter $\mathcal{U}^{\prime}$ which follows to be unique by 5.4. Therefore, we conclude that the obtained homomorphism $\pi_{\mathcal{U}^{\prime}}\colon\Lambda\to\operatorname{Sym}(f_{*}(\mathcal{U}^{\prime}))$ satisfies that the centralizer of $\pi_{\mathcal{U}^{\prime}}(\Gamma_{\ell})$ is isomorphic to $\prod^{\mathrm{alg}}_{\mathcal{U}^{\prime}}\operatorname{Alt}(n)$. In particular, this implies that $\prod^{\mathrm{alg}}_{\mathcal{U}}\operatorname{Alt}(n)$ is isomorphic to $\prod^{\mathrm{alg}}_{\mathcal{U}^{\prime}}\operatorname{Alt}(n)$. This finishes the proof of the first statement; the final claim follows from Theorem 4.7. ∎
Remark 5.6.
The proof of Theorem 5.5 also shows that the metric ultraproducts ${\rm Sym}(\mathcal{U})$ for $\mathcal{U}\in\mathbb{U}$ are pairwise non-elementarily equivalent in the sense of FO-model theory of metric structures. Indeed, if $\gamma\neq\gamma^{\prime}\colon\mathbb{P}_{\geq 7}\to\{0,1\}$, then there exists a prime $q\in\mathbb{P}_{\geq 7}$, such that $\gamma(q)\neq\gamma^{\prime}(q)$. Then, the sentence $\varphi(a_{q,\gamma(q)},q)$ distinguishes the centralizers of all homomorphisms from $\Lambda$ to the respective universal sofic groups, satisfying Conditions (1) and (2) in the proof of Theorem 5.5. Since we did not introduce the language of FO-model theory of metric structures, we omit the details.
Remark 5.7.
It is a well-known open problem in the realm of operator algebras, first formulated by Sorin Popa, to decide if there are uncountably many non-isomorphic tracial ultraproducts of matrix algebras. This question has been answered assuming the negation of the continuum hypothesis (see [farah]), but there is no unconditional result. Each tracial ultraproduct of matrix algebras contains a canonical diagonal subalgebra, whose Weyl group, i.e. the quotient of its normalizer by its centralizer, can be shown to be isomorphic to the corresponding universal sofic group. Thus, as a consequence of our results, it follows that for pairs of ultrafilters from $\mathbb{U}$, the corresponding tracial ultraproducts of matrix algebras cannot be isomorphic in a way that respects the diagonal subalgebras.
We were unable to show that the diagonal subalgebra is automatically preserved in a suitable sense. However, even though this diagonal subalgebra in not a Cartan subalgebra [popa83, popa1], the size of its normalizer should make the situation very special.
Acknowledgments
The first author acknowledges funding by the Deutsche Forschungsgemeinschaft (SPP 2026 “Geometry at infinity”). We thank Ilijas Farah for helpful comments, corrections and remarks that helped to improve the exposition.